# Equal contribution. * Corresponding author

### v1

@@ Abstract
In this work, we present ThermoHands, a new benchmark for thermal image-based egocentric 3D hand pose estimation, aimed at overcoming challenges like varying lighting and obstructions (e.g., handwear). The benchmark includes a diverse dataset from 28 subjects performing hand-object and hand-virtual interactions, accurately annotated with 3D hand poses through an automated process. We introduce a bespoken baseline method, TheFormer, utilizing dual transformer modules for effective egocentric 3D hand pose estimation in thermal imagery. Our experimental results highlight TheFormer’s leading performance and affirm thermal imaging’s effectiveness in enabling robust 3D hand pose estimation in adverse conditions.
@@ 1 Introduction
In contrast to NIR-based methods, thermal imaging cameras offer a passive sensing solution for hand pose estimation by capturing long-wave infrared (LWIR) radiation emitted from objects, thereby eliminating reliance on the visible light spectrum [54]. This unique attribute of thermal imaging introduces several benefits for 3D hand pose estimation. Primarily, it accentuates the hand’s structure via temperature differentials, negating the effects of lighting variability. Moreover, thermal cameras are capable of detecting hands even under handwear such as gloves by identifying heat transmission patterns. This ability ensures a stable and consistent representation of hands, independent of any coverings, thereby broadening the scope and reliability of hand pose estimation across various scenarios.
@@
Building on the above insights, this study probes the following research question: Can egocentric thermal imagery be effectively used for 3D hand pose estimation under various conditions (such as different lighting and handwear), and how does it compare to techniques using RGB, NIR, and depth [^1] For readability, we treat depth and NIR as two ‘spectra’, despite their usual overlap. spectral imagery? To answer this, we introduce ThermoHands, the first benchmark specifically tailored for egocentric 3D hand pose estimation utilizing thermal imaging. This benchmark is supported by a novel multi-spectral and multi-view dataset designed for egocentric 3D hand pose estimation and is unique in comprising thermal, NIR, depth, and RGB images (cf. Tab. 1). Our dataset emulates real-world application contexts by incorporating both hand-object and hand-virtual interaction activities, with participation from 28 subjects to ensure a broad representation of actions (cf. Fig. 1). To offer a thorough comparison across spectral types, we gather data under five distinct scenarios, each characterized by varying environments, handwear, and lighting conditions (cf. Tab. 2). Considering the challenges associated with manually annotating large-scale 3D hand poses, we developed an automated annotation pipeline. This pipeline leverages multi-view RGB and depth imagery to accurately and efficiently generate 3D hand pose ground truths through optimization based on the MANO model [68] (cf. Fig. 3).
@@
Together with the multi-spectral dataset, we introduce a new baseline method named TheFormer, specifically designed for thermal image-based egocentric 3D hand pose estimation (cf. Fig. 4). This approach is notable for its two parallel transformer modules, i.e., mask-guided spatial transformer and temporal transformer, which encode spatio-temporal relationship for 3D hand joints without losing the computation efficiency. Our validation process begins with verifying the annotation quality, which averages an accuracy of 1cm (cf. Tab. 3). We then benchmark TheFormer against leading methods (cf. Tab. 4) and compare the performance of various spectral images (cf. Tab. 5 and Fig. 6). The findings underscore thermal imagery’s advantages in difficult lighting conditions and when hands are gloved, showing superior performance and better adaptability to challenging settings than other spectral techniques. Our main contributions are summarized as follows:
@@
    We introduce a new baseline method, termed TheFormer, and implement two state-of-the-art image-based methods on our dataset for benchmarking.
@@
    Based on the ThermoHands benchmark, we conduct comprehensive experiments and analysis on TheFormer and state-of-the-art methods.
@@ 2 Related Works
Despite the existing progress, previous datasets only provide depth [97, 62], RGB images [61, 60, 104, 64] or both of them [23, 85, 29, 103, 63, 83, 11, 52, 27, 7, 45] as the input spectra, unable to support the study of NIR or thermal image-based 3D hand pose estimation. ThermoHands fills the gap by providing a moderate amount of multi-spectral image data, from infrared to visual light, paired with depth images. Moreover, we capture bimanual actions from both egocentric and exocentric viewpoints and design hand-object as well as hand-virtual interaction actions to facilitate a wide range of applications. Tab. 1 shows a comprehensive comparison between the existing datasets (in chronological order) and ours.
Table 1: Comparison of ThermoHands with existing image-based hand datasets with 3D pose annotations. (*) The number of synchronized frames of images. (**) No markers or sensors are attached to hands during data creation or capture.
Dataset (*)frames (**)markless depth thermal real ego exo two-hand hand-obj int hand-virt int
BigHands2.2M [97] 2.2M
SynthHands [63] 220K
EgoDexter [63] 3K
FPHA [23] 105K
ObMan [29] 150K
FreiHAND [104] 37K
HO-3D [27] 78K
ContactPose [7] 2.9M
InterHand2.6M [60] 2.6M
H2O [45] 571K
DexYCB [11] 508K
HOI4D [52] 2.4M
AssemblyHands [64] 3.0M
ARCTIC [18] 2.1M
ThermoHands (ours) 96K
@@ 3.1 Multi-Spectral Hand Pose Dataset
Overview. At the core of our benchmark lies a multi-spectral dataset for 3D hand pose estimation (cf. Tab. 2), capturing hand actions performed by 28 subjects of various ethnicities and genders [^2] The study has received the ethical approval from XXX (no name for blind review), and participant consent forms were signed before the collection.. As shown in Fig. 1, we develop a customized head-mounted sensor platform (HMSP) and an exocentric platform to record multi-view data. During capture, our participants are asked to perform pre-defined hand-object and hand-virtual interaction actions within the playground above the table. The main part is captured in the well-illuminated office scenario. To facilitate the evaluation under different settings, four auxiliary parts are recorded i) under the darkness, ii) under the sun glare, iii) with gloves on hand, and iv) in the kitchen environment with different actions, respectively.
Sensor Platforms. The HMSP is mounted with an Intel RealSense L515 LiDAR depth camera [33] streaming egocentric RGB, depth and NIR images, and a FLIR Boson 640 thermal camera [87] receiving the LWIR to obtain the thermal images. To better simulate XR devices, the HMSP is modularly designed as two components: a base and a sensor board with a fixed 30-degree downward tilt. An extra exocentric platform equipped with an Intel RealSense D455 [32] is leveraged to support multi-view annotation (cf. Sec. 3.2) as well as provide the RGB-D image data from the third-person viewpoint. As exhibited in Fig. 1, we place the two depth sensors outside each other’s FoV to minimize interference caused by their NIR emitters [33, 32]. In the office environment, random heat sources, e.g., servers and chargers, are strategically placed in the background to increase realism and introduce challenging factors into thermal images.
Synchronization. We use a single PC to simultaneously gather data streams from two sensor platforms, ensuring the synchronization of their timestamps. After collection, we synchronize six types of images, each with distinct frame rates, w.r.t. the timestamps of thermal images (8.5fps), thereby generating synchronized multi-spectral, multi-view data samples as our released data.
Well-illuminated office (Main) Other settings Total
@@
Calibration. For accuracy, we use factory-calibrated intrinsic parameters from the D455 and L515 cameras [33, 32]. To better calibrate the parameters of the thermal camera, we self-design a modular calibration chessboard as shown in Fig. 2. Before calibration, we cool down the black base board while keeping the white cubes at room temperature to create a visible chessboard pattern in all three spectra, which can be seen in Fig. 2. In this way, we can simultaneously calibrate the intrinsic parameter of the thermal camera and its extrinsic parameter w.r.t. the L515 camera. To enable the calibration between the two viewpoints, we place 11 IR reflective spheres at random locations and heights above the table. At the first frame of each sequence, we manually annotate their 2D locations from two viewpoints, retrieve their depth values and compute the transformation between two viewpoints by solving the PnP [19]. For the subsequent frames, the KISS-ICP [90] odometry method is used to track the motion of the ego-head using the point clouds converted from the egocentric depth images.
@@
Dataset Statistics. As shown in Tab. 2, our dataset consists of approximately 96K synchronized multi-spectral multi-view frames (cf. Fig. 1) and 352 independent sequences in total. The main part is collected under the well-illuminated office scenario, where each participant [^3] Due to their limited time, 7 participants only perform the hand-object actions. performs 7 scenario-specific hand-object interaction actions: cut paper, fold paper, pour water, read book, staple paper, write with pen, write with pencil, and 5 hand-virtual interaction actions: pinch and drag, pinch and hold, swipe, tap, touch, with two hands. This main part is divided into the training, validation and testing splits by subjects with a ratio of 4:1:2. We also collect four auxiliary testing sets by asking one subject to perform the aforementioned 12 actions in the darkness, sun glare and gloves settings individually, and two subjects to perform 7 scenarios-specific interaction actions: cut, spray, stir, wash hands, wash mug, wash plate, wipe in the kitchen environment. Please refer to the supplementary for more dataset details.
Figure 3: Automatic annotation pipeline of 3D hand pose. We utilize the multi-view RGB and depth images as the input source and retrieve both 2D and 3D constraint information. Various error terms are formulated to optimize the MANO parameters.
@@ 3.2 Hand Pose Annotation
To refrain from employing tedious human efforts for annotation, we implement a fully automatic annotation pipeline, similar to the approaches in [27, 7, 45], to obtain the 3D hand pose ground truth for our dataset.
In particular, we use the MANO statistical hand model [68] to represent 3D hand pose. The MANO model parameterizes the hand mesh vertices $\mathcal{V}(\beta,\theta)$ into two low-dimensional embeddings, i.e., the shape parameters $\beta\in\mathbb{R}^{10}$ and the pose parameters $\theta\in\mathbb{R}^{51}$, consisting of 45 parameters accounting for 15 hand joint angles (3 DoF for each) plus the rest for global rotation and translation. Following the MANO PyTorch version in [30], we denote the $N_{\mathcal{J}}=21$ hand joints mapped from the hand parameters as $\mathcal{J}(\beta,\theta)$. The MANO fitting is performed by minimizing the following optimization objective per frame for each hand:
@@
where $\lambda_{j2d},\lambda_{mask},\lambda_{j3d},\lambda_{mesh},\lambda_{reg}$ are used to balance the weight of different errors. The diagram illustrating our annotation process is shown in Fig. 3.
Initialization. For each sequence, we optimize the shape parameter $\beta$ [^4] For simplicity, we omit $\beta$ in Eq. 1 and subsequent equations. together with $\theta$ only at the first frame using Eq. 1 until convergence and fix its values for the subsequent frames. We initialize the pose parameter $\theta$ for each frame using the optimization result from the last frame. This helps to accelerate the convergence as well as keep the temporal consistency.
2D Joint Error $\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}$. Given RGB images from $N_{C}$ viewpoints, we infer 2D hand joints $\mathcal{J}^{2D}$ using MediaPipe Hands [53] and define the 2D joint error as:
$\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}=\sum_{c=1}^{N_{C}}\alpha_{c}\sum_{i=1}^{N_{\mathcal{J}}}||\mathcal{J}^{2D}_{c,i}-\pi_{c}(\mathcal{J}(\theta)_{i})||$ (2)
where $\pi_{c}(\cdot)$ returns the 2D projection location for 3D position in the $c$-th camera viewpoint, and $\alpha_{c}$ is hyperparameter used to weigh different viewpoints.
2D Mask Error $\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}$. To generate the high-quality 2D hand mask, we prompt the prevalent Segment Anything Model [42] with the 2D hand joints $\mathcal{J}^{2D}$ and the bounding box derived from it. We penalize the distance between the hand mesh vertices $\mathcal{V}(\theta)$ and the 2D binary hand mask $\mathcal{M}_{c}$ as:
$\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}=\sum_{c=1}^{N_{C}}\alpha_{c}\sum_{i=1}^{N_{\mathcal{V}}}\min_{j}||\mathcal{M}_{c,j}-\pi_{c}(\mathcal{V}(\theta)_{i})||$ (3)
where $\mathcal{M}_{c,j}$ is the coordinate of $j$-th non-zero pixel in the mask $\mathcal{M}_{c}$.
3D Joint Error $\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}$. We triangulate the 2D joints from multiple views to lift them to 3D joints $\mathcal{J}^{3D}$ and measure their difference to $\mathcal{J}(\theta)$, which is written as:
$\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}=\sum_{i=1}^{N_{\mathcal{J}}}||\mathcal{J}^{3D}_{i}-\mathcal{J}(\theta)_{i}||$ (4)
3D Mesh Error $\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}$. To better supervise the hand mesh and fasten the optimization, we generate the 3D hand pose cloud $\mathcal{P}$ by cropping the depth image using the 2D hand mask and merging all views together. The 3D mesh error term compensates for the distance between the hand mesh and point cloud:
$\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}=\sum_{i=1}^{N_{\mathcal{V}}}\min_{j}||\mathcal{P}_{j}-\mathcal{V}(\theta)_{i}||$ (5)
Regularization $\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}$. To alleviate irregular hand articulation, we constrain the joint angles to pre-defined lower and upper boundaries $\underline{\theta}$ and $\overline{\theta}$:
$\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}=\sum_{i=1}^{45}(\max(\underline{\theta_{i}}-\theta_{i},0)+\max(\theta_{i}-\overline{\theta_{i}},0))$ (6)
Note that we also impose regularization to the shape parameter $\beta$ during initialization. Please see details of our annotation in supplementary material.
@@
Figure 4: Overall Framework of TheFormer. Given thermal images, we extract both low-level and high-level features for each frame, which are input to the mask-guided spatial transformer and temporal transformer. Frame-wise spatial and temporal embeddings are concatenated and fed into the pose head to regress the 3D hand pose.
3.3 TheFormer: A Baseline Method
Our proposed baseline method, dubbed TheFormer, learns 3D hand pose estimation from egocentric thermal images based on Transformer, as exhibited in Fig. 4. The baseline network features in its two parallel transformer modules, i.e., mask-guided spatial transformer and temporal transformer, to model the spatio-temporal relationship of hand joints while being computationally efficient.
@@ 3.3 TheFormer: A Baseline Method
Problem Definition. We consider two problem settings: single image-based and video-based egocentric 3D hand pose estimation for our benchmark. In the former setting, we aim to estimate the 3D joint positions $\mathcal{J}_{t}$ for two hands given the single thermal image $\mathcal{I}_{t}$ captured for the $t$-th frame. For the video-based one, our input is a sequence of thermal images $\mathcal{S}=\{\mathcal{I}_{i}\}_{i=1}^{T}$ and we estimate the per-frame 3D hand joint positions $\mathcal{J}_{i}$ together. Compared to the single image-based counterpart, the video-based setting allows for a fully temporal interaction among sequential images, potentially offering a better performance. Nevertheless, the accumulation of sequential images leads to a higher latency for online inference, limiting its usage to applications that demand real-time hand tracking.
Network Architecture. Without losing the generality, here we illustrate our network architecture for the video-based setting. Note that our network can be flexibly adapted to the single image-based problem by setting $T=1$. As seen in Fig. 4, given sequential thermal images, our architecture extracts the multi-level 2D feature using the backbone network per frame. We design a mask-guided spatial transformer module to enhance the spatial representation while enforcing the network to focus on the hand area. In parallel, we apply a temporal transformer module to reason the feature interaction along the temporal dimension. Both two transformers output per-frame feature embeddings that encode the spatial and temporal information, respectively. We concatenate them and use a pose head to regress the 3D coordinates of hand joints for two hands.
Backbone Features. For efficiency, we use the lightweight ResNet-18 [31] as our backbone network to extract multi-level features from thermal images. Specifically, we reserve both the low-level and high-level image features and input them to the spatial and temporal transformer modules individually. Low-level features carry fine-grained spatial details, which is crucial for accurate mask generation and spatial reasoning in our mask-guided spatial transformer. High-level features provide a broader context, helping temporal transformers to better understand the overall scene dynamics and the hand’s role within it.
@@
Mask-guided Spatial Transformer. Human hands are highly articulated objects that can adopt a wide variety of poses, often against complex backgrounds. We propose a mask-guided spatial transformer module to accurately identify and focus on the intricacies of hand poses during spatial feature interaction. Given low-level features, we first utilize a mask head to estimate the binary hand mask in the thermal image. Then, we leverage the deformable self-attention [102] to refine the hand spatial features under the guidance of the hand mask. Specifically, we only take feature elements whose spatial locations are within the hand area as queries and sample keys from only the hand area and its surrounding locations. In this way, we not only reduce the computation waste on the irrelevant region but also increase the robustness to background clutter. Lastly, we reduce the spatial dimensions of the spatial features via several downsample convolutional layers and flatten it into 1D spatial feature embedding per frame.
@@
Temporal Transformer. Temporal information is crucial for 3D hand pose estimation when coping with occlusion and solving ambiguities. To efficiently model temporal relationships, we apply the average-pooling to the high-level features to get the global feature vector per frame. The self-attention [88] is then employed to explicitly attend to the feature vector of every frame for a comprehensive temporal interaction. Similar to the spatial transformer, the output is frame-wise feature embedding that encompasses the temporal information.
Pose Regression. We concatenate the frame-wise spatial and temporal embedding from two transformer modules and get the spatio-temporal representations for each frame. In the pose head, we use the MLP to project the representations to the output space and obtain the per-frame 3D joint output $\mathcal{J}_{i}$.
@@
Training Loss. We leverage the binary cross-entropy loss to supervise the hand segmentation with the mask ground truth rendered from the annotated hand mesh (cf. Sec. 3.2). For 3D hand joint positions, we measure the L1 distance of its 2D projection and depth to that of the ground truth separately.
For the video-based setting, our baseline method is named as TheFormer-V, which accepts sequential images as input. The variant of our baseline to the single image-based setting is called TheFormer-S, where the temporal transformer degrades to a global transformer that only encodes the global information for a single frame. Please see the supplementary materials for more parameters, design and training details of our baseline method.

### v3

# ThermoHands: A Benchmark for 3D Hand Pose Estimation from Egocentric Thermal Images

@@ Abstract
In this work, we present ThermoHands, a new benchmark for thermal image-based egocentric 3D hand pose estimation, aimed at overcoming challenges like varying lighting conditions and obstructions (e.g., handwear). The benchmark includes a multi-view and multi-spectral dataset collected from 28 subjects performing hand-object and hand-virtual interactions under diverse scenarios, accurately annotated with 3D hand poses through an automated process. We introduce a new baseline method, TherFormer, utilizing dual transformer modules for effective egocentric 3D hand pose estimation in thermal imagery. Our experimental results highlight TherFormer’s leading performance and affirm thermal imaging’s effectiveness in enabling robust 3D hand pose estimation in adverse conditions.
@@ 1 Introduction
In contrast to NIR-based methods, thermal imaging cameras offer a passive sensing solution for hand pose estimation by capturing long-wave infrared (LWIR) radiation emitted from objects, thereby eliminating reliance on the visible light spectrum [24]. This unique attribute of thermal imaging introduces several benefits for hand pose estimation. Primarily, it accentuates the hand’s structure via temperature differentials, negating the effects of lighting variability. Moreover, thermal cameras can detect hands even under handwear such as gloves by identifying heat transmission patterns. This ability ensures a stable and consistent representation of hands, independent of any coverings, thereby broadening the scope and reliability of hand pose estimation across various scenarios.
@@
Building on the above insights, this study probes the following research question: Can egocentric thermal imagery be effectively used for 3D hand pose estimation under various conditions (such as different lighting and handwear), and how does it compare to techniques using RGB, NIR, and depth [^1] For readability, we treat depth and NIR as two ‘spectra’, despite their usual overlap. spectral imagery? To answer this, we introduce ThermoHands, the first benchmark specifically tailored for egocentric 3D hand pose estimation utilizing thermal imaging. This benchmark is supported by a novel multi-spectral and multi-view dataset designed for egocentric 3D hand pose estimation and is unique in comprising thermal, NIR, depth, and RGB images (cfFig. 1). Our dataset emulates real-world application contexts by incorporating both hand-object and hand-virtual interaction activities, with participation from 28 subjects to ensure a broad representation of actions. To offer a thorough comparison across spectral types, we gather data under five distinct scenarios, each characterized by varying environments, handwear, and lighting conditions (cfTab. 1). Considering the challenges associated with manually annotating large-scale 3D hand poses, we developed an automated annotation pipeline. This pipeline leverages multi-view RGB and depth imagery to accurately and efficiently generate 3D hand pose ground truths through optimization based on the MANO model [25] (cfFig. 3).
@@
Together with the multi-spectral dataset, we introduce a new baseline method named TherFormer, specifically designed for thermal image-based egocentric 3D hand pose estimation (cfFig. 4). This approach is notable for its two consecutive transformer modules, i.e., mask-guided spatial transformer and temporal transformer, which encode spatio-temporal relationship for 3D hand joints without losing the computation efficiency. Our validation process begins with verifying the annotation quality, which averages an accuracy of 1cm (cfTab. 2). We then benchmark TherFormer against leading methods (cfTab. 3) and compare the performance of various spectral images (cfTab. 4 and Fig. 6). The findings underscore thermal imagery’s advantages in difficult lighting conditions and when hands are gloved, showing superior performance and better adaptability to challenging settings than other spectral techniques. Our main contributions are summarized as follows:
@@
    We introduce a new baseline method, termed TherFormer, and implement state-of-the-art image-based methods on our dataset for benchmarking.
@@
    Based on the ThermoHands benchmark, we conduct comprehensive experiments and analysis on TherFormer and state-of-the-art methods.
@@ 2 Related Works
Despite the existing progress, previous datasets only provide depth [26, 32], RGB images [33, 38, 37, 40] or both of them [27, 28, 31, 34, 30, 35, 36, 39, 41, 42, 43] as the input spectra, unable to support the study of NIR or thermal image-based 3D hand pose estimation. ThermoHands fills the gap by providing a moderate amount of multi-spectral image data, from infrared to visual light, paired with depth images. Moreover, we capture bimanual actions from both egocentric and exocentric viewpoints and design hand-object as well as hand-virtual interaction actions to facilitate a wide range of applications.
@@ 3.1 Multi-Spectral Hand Pose Dataset
Overview. At the core of our benchmark lies a multi-spectral dataset for 3D hand pose estimation (cfTab. 1), capturing hand actions performed by 28 subjects of various ethnicities and genders [^2] The study has received the ethical approval from School of Informatics, University of Edinburgh, and participant consent forms were signed before the collection.. As shown in Fig. 1, we develop a customized head-mounted sensor platform (HMSP) and an exocentric platform to record multi-view data. During the capture, our participants are asked to perform pre-defined hand-object and hand-virtual interaction actions within the playground above the table. The main part is captured in the normal office scenario. To facilitate the evaluation under different settings, four auxiliary parts are recorded i) under the darkness, ii) under the sun glare, iii) with gloves on hand, and iv) in the kitchen environment with different actions, respectively.
Normal office (Main) Other settings Total
@@
Sensor Platforms. The HMSP is mounted with an Intel RealSense L515 LiDAR depth camera [20] streaming egocentric RGB, depth and NIR images, and a FLIR Boson 640 thermal camera [96] receiving the LWIR to obtain the thermal images. To better simulate XR devices, the HMSP is modularly designed as two components: a base and a sensor board with a fixed 30-degree downward tilt. An extra exocentric platform equipped with an Intel RealSense D455 [21] is leveraged to support multi-view annotation (cf. Sec. 3.2) as well as provide the RGB-D image data from the third-person viewpoint. As exhibited in Fig. 1, we place the two depth sensors outside each other’s field of view (FoV) to minimize interference caused by their NIR emitters [20, 21]. In the office environment, random heat sources, e.g., servers and chargers, are strategically placed in the background to increase realism and introduce challenging factors into thermal images.
Synchronization. We use a single PC to simultaneously gather data streams from two sensor platforms, ensuring the synchronization of their timestamps. After collection, we synchronize six types of images, each with distinct frame rates, w.r.t. the timestamps of thermal images (8.5fps), thereby generating well-aligned multi-spectral, multi-view data samples as our released data.
@@
Calibration. For accuracy, we use factory-calibrated intrinsic parameters from the D455 and L515 cameras [20, 21]. To better calibrate the parameters of the thermal camera, we self-design a modular calibration chessboard as shown in Fig. 2. Before calibration, we cool down the black base board while keeping the white cubes at room temperature to create a visible chessboard pattern in all three spectra, which can be seen in Fig. 2. In this way, we can simultaneously calibrate the intrinsic parameter of the L515 thermal camera and its extrinsic parameter w.r.t. the D455 RGB-D camera. To enable the calibration between the two viewpoints, we place 11 IR reflective spheres at random locations and heights above the table. At the first frame of each sequence, we manually annotate their 2D locations from two viewpoints, retrieve their depth values and compute the transformation between two viewpoints by solving the PnP problem [97]. For the subsequent frames, as the exocentric platform keeps stationary, we only track the the motion of the ego-head using KISS-ICP [98] odometry method, which takes the point clouds converted from the egocentric depth image as input.
@@
Dataset Statistics. As seen in Tab. 1, our dataset consists of approximately 96K synchronized multi-spectral multi-view frames (cfFig. 1) and 352 independent sequences in total. The main part is collected under the normal office scenario, where each participant [^3] Due to their limited time, 7 participants only perform the hand-object actions. performs 7 hand-object interaction actions: cut paper, fold paper, pour water, read book, staple paper, write with pen, write with pencil, and 5 hand-virtual interaction actions: pinch and drag, pinch and hold, swipe, tap, touch, with two hands. This main part is divided into the training, validation and testing splits by subjects with a ratio of 4:1:2. We also collect four auxiliary testing sets by asking one subject to perform the aforementioned 12 actions in the darkness, sun glare and gloves settings individually, and two subjects to perform 7 scenarios-specific interaction actions: cut, spray, stir, wash hands, wash mug, wash plate, wipe in the kitchen. Please refer to the supplementary for more dataset details.
@@ 3.2 Hand Pose Annotation
To avoid employing tedious human efforts for annotation, we implement a fully automatic annotation pipeline, similar to the approaches in [41, 42, 43], to obtain the 3D hand pose ground truth for our dataset. In particular, we use the MANO statistical hand model [25] to represent 3D hand pose. The MANO model parameterizes the hand mesh vertices into two low-dimensional embeddings, i.e., the shape parameters $\beta\in\mathbb{R}^{10}$ and the pose parameters $\theta\in\mathbb{R}^{51}$. The MANO fitting is performed by minimizing the following optimization objective per frame for each hand:
@@
where $\lambda_{j2d},\lambda_{mask},\lambda_{j3d},\lambda_{mesh},\lambda_{reg}$ are used to balance the weight of different errors. The diagram illustrating our annotation process is shown in Fig. 3. For more explanation of error terms and implementation details of our dataset annotation process, please refer to our supplementary document.
Figure 3: Automatic annotation pipeline of 3D hand pose. We utilize the multi-view RGB and depth images as the input source and retrieve constraint information with off-the-shelf MediaPipe Hands [99] and SAM [100]. Various error terms are formulated to optimize the MANO parameters.
3.3 TherFormer: A Baseline Method
For our benchmark, we setup a baseline method, dubbed TherFormer, for thermal image-based 3D hand pose estimation. As exhibited in Fig. 4, TherFormer features in its two consecuative transformer modules to model the spatio-temporal relationship of hand joints while being computationally efficient. Note that TherFormer is also capable of processing other spectral images due to their format consistency with thermal images. However, our primary objective is to establish a baseline method for future research, rather than focusing exclusively on methods for thermal images.
@@ 3.3 TherFormer: A Baseline Method
Figure 4: Overall Framework of TherFormer. Backbone features are input to the mask-guided spatial transformer and temporal transformer to enhance the spatial representation and temporal interaction. Spatio-temporal embeddings are fed into the pose head to regress the 3D hand pose.
@@
Problem Definition. We consider two problem settings: single image-based and video-based egocentric 3D hand pose estimation for our benchmark. In the former setting, we aim to estimate the 3D joint positions $\mathcal{J}_{t}$ for two hands given the single thermal image $\mathcal{I}_{t}$ captured for the $t$-th frame. For the video-based one, our input is a sequence of thermal images $\mathcal{S}=\{\mathcal{I}_{i}\}_{i=1}^{T}$ and we estimate the per-frame 3D hand joint positions $\mathcal{J}_{i}$ together. Without losing the generality, here we illustrate our network architecture for the video-based setting, i.e., TherFormer-V. In practice, our network can be flexibly adapted to the single image-based version (i.e., TherFormer-S) by setting $T=1$.
@@
Mask-guided Spatial Transformer. Human hands are highly articulated objects that can adopt a wide variety of poses, often against complex backgrounds. We propose a mask-guided spatial transformer module to accurately identify and focus on the intricacies of hand poses during spatial feature interaction. Given backbone features, we first utilize a mask head to estimate the binary hand mask in the thermal image. Then, we leverage the deformable self-attention [101] to refine the hand spatial features under the guidance of the estimated hand mask. Specifically, we only take feature elements whose spatial locations are within the hand area as queries and sample keys from only the hand area and its surrounding locations. In this way, we not only reduce the computation waste on the irrelevant region but also increase the robustness to background clutter. Lastly, we reduce the spatial dimensions of the spatial features with a series of convolutions layers for efficiency.
@@
Temporal Transformer. Temporal information is crucial for 3D hand pose estimation when coping with occlusion and solving ambiguities. To model temporal relationships, we first flatten the spatial features into 1D feature vectors and then employed the temporal self-attention [102] to explicitly attend to the feature vector of every frame. The output is frame-wise spatio-temporal feature embeddings. Note that the temporal self-attention degrades to an MLP for single-image based setting.
@@
In the pose head, we use the MLP to project the embeddings to the output space and obtain the per-frame 3D joint $\mathcal{J}_{i}$. We leverage the binary cross-entropy loss to supervise the hand segmentation with the mask ground truth rendered from the annotated hand mesh (cfSec. 3.2). For 3D hand joint positions, we measure the L1 distance of its 2D projection and depth to that of the ground truth separately. Please see more details of TherFormer in our supplementary document.